# SDSS-IV MaNGA: how do star-formation histories affect gas-phase abundances?

### 11 Dec 2023 v1

@@ Abstract
On galaxy scales, both O/H and N/O have been found to correlate significantly with stellar mass (e.g. Tremonti et al., 2004; Pérez-Montero et al., 2013; Masters et al., 2016), with the former relation typically referred to as the mass–metallicity relation (MZR). The MZR has been found to display a residual dependence on star-formation rate (SFR), with higher SFRs associated with lower metallicities at a given mass, producing a three-way relation known as the fundamental metallicity relation (FMR; e.g. Ellison et al., 2008; Lara-López et al., 2010; Mannucci et al., 2010). On the other hand, the FMR’s reduction in scatter from the MZR appears to be relatively modest and is seen mostly in lower-mass () galaxies (e.g. Salim et al., 2014), which are potential factors in the reported non-detections of a galaxy-scale FMR in smaller IFU data-sets (e.g. Sánchez et al., 2013, 2017b; Barrera-Ballesteros et al., 2017; Alvarez-Hurtado et al., 2022). Cresci et al. (2019) have investigated this point for CALIFA and MaNGA data, and they argue that a galaxy-scale FMR remains apparent when the relation’s mass dependence is considered.
@@ 2.3 Gas-phase abundance calculation
A range of gas abundance calibrators have been proposed in the literature, in which gas metallicities – along with N/O abundance ratios – are estimated from emission line flux ratios. Many calibrators have been derived from photoionization models, in which input chemical abundances allow predicted emission line ratios to be recovered (e.g. Blanc et al., 2015; Dopita et al., 2016; Morisset et al., 2016). However, such calibrators depend on major assumptions concerning the physical behaviours of ionizing populations’ atmospheres; moreover, such calibrators assume O/H to trend with ionisation strength or else with abundance ratios such as N/O, implicitly or otherwise (for a summary, see Section 3.1 of Sánchez et al., 2017b, and references therein).
@@ 3 Results
Considering the Spearman rank correlation coefficient , with find D4000 and to both correlate slightly stronger with N/O than with the gas metallicity. For D4000, we find and for N/O and for metallicity respectively; we find in both cases and in all subsequent cases, where describes the probability of obtaining coefficients equal to or higher than what is measured in the case where there is no underlying correlation. For , we find and for N/O and for metallicity respectively. In the case of , we find = 0.59 and = 0.61 for N/O and for metallicity respectively, meaning what we obtain a slightly higher correlation with metallicity. As such, we find both gas abundance parameters to correlate strongly with properties relating to the stellar components, as has been reported repeatedly for the parameters individually in MaNGA and elsewhere (e.g. Sánchez et al., 2013; Barrera-Ballesteros et al., 2016; Sánchez et al., 2017a; Hwang et al., 2019; Espinosa-Ponce et al., 2022; Schaefer et al., 2022); however, neither parameter is clearly more or less informative than the other. We also note that excluding GV subsample spaxels has very little impact on these comparisons, due to the relatively small size of the GV subsample.
@@ 4 Discussion
To further assess the behaviour of starburst and post-starburst spaxels, we compare in Figure 8 the starburst sample to a new ‘control sample’ selected from post-starburst samples. This control sample was selected using a combination of and , just as for the control samples already discussed. In this case, we find the starburst and ‘control’ PSB sample to differ in two key ways: the metallicity and N/O of starbursts is lower on average than for PSB regions, while starburst spaxels’ metallicity is somewhat elevated compared to PSB spaxels at the highest N/O values. Our findings therefore support a view in which gas-phase abundances typically increase rapidly after a starburst, such that post-starburst spaxels behave very similarly to star-forming spaxels in terms of their chemical abundances. A rapid enrichment of gas is similar to what Leung et al.(in prep) report for MaNGA post-starburst regions’ stellar populations. Leung et al.(in prep) obtain the metallicities of old and young populations in these regions separately by performing full spectral fitting with the Bagpipes code (Carnall et al., 2018; Carnall et al., 2019), and they find younger populations – those formed during the starburst – to typically be more metal-rich than their older counterparts at a level reasonably consistent with the metallicity offset reported between star-forming and quiescent galaxy samples (e.g. Peng et al., 2015). It should be noted that a rise in metallicity over successive stellar generations is entirely expected, and it has been reported previously from fossil record analyses of spectra (e.g. Panter et al., 2008).
@@ 5 Summary & Conclusion
We found that starburst regions possess diluted abundances when compared to ordinary star-forming regions, consistent with previous work. We also found post-starburst regions to behave very similarly to star-forming regions in N/O–O/H space, supporting a view in which gas is enriched rapidly following a starburst event; such a view is consistent with the stellar populations observed within MaNGA post-starburst regions, in which young stars are typically found to be significantly more metal-rich than older stars (Leung et al. in prep). Furthermore, we found green-valley regions to posses elevated N/O abundances on average for their metallicity, which appears to relate to the behaviour of high-D4000 galaxy regions in our sample more generally. We argued that such regions may have experienced recent metal-poor inflow and that they were possibly rejuvenated as a consequence, with SFE variations being another possible factor. We also found various stellar population parameters (, , D4000) along with galactocentric radius to trend strongly with both O/H and N/O, with neither gas abundance parameter being obviously more or less informative than the other for tracing the evolution of galaxy regions.

### 19 Jan 2024 v2

@@ Abstract
On galaxy scales, both O/H and N/O have been found to correlate significantly with stellar mass (e.g. Tremonti et al., 2004; Pérez-Montero et al., 2013; Masters et al., 2016), with the former relation typically referred to as the mass–metallicity relation (MZR). The MZR has been found to display a residual dependence on star-formation rate (SFR), with higher SFRs associated with lower metallicities at a given mass, producing a three-way relation known as the fundamental metallicity relation (FMR; e.g. Ellison et al., 2008; Lara-López et al., 2010; Mannucci et al., 2010). On the other hand, the FMR’s reduction in scatter from the MZR appears to be relatively modest and is seen mostly in lower-mass () galaxies (e.g. Salim et al., 2014), which are potential factors in the reported non-detections of a galaxy-scale FMR in smaller IFU data-sets (e.g. Sánchez et al., 2013, 2017; Barrera-Ballesteros et al., 2017; Alvarez-Hurtado et al., 2022). Cresci et al. (2019) have investigated this point for CALIFA and MaNGA data, and they argue that a galaxy-scale FMR remains apparent when the relation’s mass dependence is considered.
@@ 2.3 Gas-phase abundance calculation
A range of gas abundance calibrators have been proposed in the literature, in which gas metallicities – along with N/O abundance ratios – are estimated from emission line flux ratios. Many calibrators have been derived from photoionization models, in which input chemical abundances allow predicted emission line ratios to be recovered (e.g. Blanc et al., 2015; Dopita et al., 2016; Morisset et al., 2016). However, such calibrators depend on major assumptions concerning the physical behaviours of ionizing populations’ atmospheres; moreover, such calibrators assume O/H to trend with ionisation strength or else with abundance ratios such as N/O, implicitly or otherwise (for a summary, see Section 3.1 of Sánchez et al., 2017, and references therein).
@@ 3 Results
Considering the Spearman rank correlation coefficient , with find D4000 and to both correlate slightly stronger with N/O than with the gas metallicity. For D4000, we find and for N/O and for metallicity respectively; we find in both cases and in all subsequent cases, where describes the probability of obtaining coefficients equal to or higher than what is measured in the case where there is no underlying correlation. For , we find and for N/O and for metallicity respectively. In the case of , we find = 0.60 and = 0.61 for N/O and for metallicity respectively, meaning what we obtain a slightly higher correlation with metallicity. As such, we find both gas abundance parameters to correlate strongly with properties relating to the stellar components, as has been reported repeatedly for the parameters individually in MaNGA and elsewhere (e.g. Sánchez et al., 2013; Barrera-Ballesteros et al., 2016; Sánchez et al., 2017; Hwang et al., 2019; Espinosa-Ponce et al., 2022; Schaefer et al., 2022); however, neither parameter is clearly more or less informative than the other. We also note that excluding GV subsample spaxels has very little impact on these comparisons, due to the relatively small size of the GV subsample.
@@ 4 Discussion
To further assess the behaviour of starburst and post-starburst spaxels, we compare in Figure 8 the starburst sample to a new ‘control sample’ selected from post-starburst samples. This control sample was selected using a combination of and , just as for the control samples already discussed. In this case, we find the starburst and ‘control’ PSB sample to differ in two key ways: the metallicity and N/O of starbursts is lower on average than for PSB regions, while starburst spaxels’ metallicity is somewhat elevated compared to PSB spaxels at the highest N/O values. Our findings therefore support a view in which gas-phase abundances typically increase rapidly after a starburst, such that post-starburst spaxels behave very similarly to star-forming spaxels in terms of their chemical abundances. A rapid enrichment of gas is similar to what Leung et al. (2023) report for MaNGA post-starburst regions’ stellar populations. Leung et al. (2023) obtain the metallicities of old and young populations in these regions separately by performing full spectral fitting with the Bagpipes code (Carnall et al., 2018; Carnall et al., 2019), and they find younger populations – those formed during the starburst – to typically be more metal-rich than their older counterparts at a level reasonably consistent with the metallicity offset reported between star-forming and quiescent galaxy samples (e.g. Peng et al., 2015). It should be noted that a rise in metallicity over successive stellar generations is entirely expected, and it has been reported previously from fossil record analyses of spectra (e.g. Panter et al., 2008).
@@ 5 Summary & Conclusion
We found that starburst regions possess diluted abundances when compared to ordinary star-forming regions, consistent with previous work. We also found post-starburst regions to behave very similarly to star-forming regions in N/O–O/H space, supporting a view in which gas is enriched rapidly following a starburst event; such a view is consistent with the stellar populations observed within MaNGA post-starburst regions, in which young stars are typically found to be significantly more metal-rich than older stars (Leung et al., 2023). Furthermore, we found green-valley regions to posses elevated N/O abundances on average for their metallicity, which appears to relate to the behaviour of high-D4000 galaxy regions in our sample more generally. We argued that such regions may have experienced recent metal-poor inflow and that they were possibly rejuvenated as a consequence, with SFE variations being another possible factor. We also found various stellar population parameters (, , D4000) along with galactocentric radius to trend strongly with both O/H and N/O, with neither gas abundance parameter being obviously more or less informative than the other for tracing the evolution of galaxy regions.